OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle if a = 3 and b = 4. Use: a2 + b2 = c2 5 Square root of 7 25 7

OpenStudy (whpalmer4):

Pythagorean theorem is \(a^2+b^2 = c^2\). We know \(a = 3\) and \(b=4\). We want to find \(c\). \[a^2+b^2=c^2\]\[\sqrt{a^2+b^2} = c\]\[c = \sqrt{(3)^2+(4)^2}\]What is the answer?

OpenStudy (anonymous):